Hello, I need to write an equation with the given roots, write the equation in the form ax^2 + bx + c = 0. The problem is 3, -4.
Answer by richwmiller(17219)

(x-3)*(x+4)=0
x^2-3x+4x-12=0
x^2+x-12=0